# Asymmetric Symmetry Breaking: Unequal Probabilities of Vacuum Selection

### v1

@@ III.2 Cosine-type potential
Figure 3: a. The cosine-type potential $V(\phi)$ (blue) in Eq.(8) with $\beta_{-}=2$ and $\beta_{+}=6$, and its derivative $V^{\prime}(\phi)$ (red); b. Time time evolution of the scalar fields. Each has three cases to approach positive and negative minimums. The inset plot shows the very early fluctuations around the top due to the external temperature. After some time, they will persistently roll down the potential as the arrows indicate; c. Relation between the probability $P_{+}$ to the ratio $\beta_{+}/\beta_{-}$. Numerical results match the theoretical prediction (red line) Eq.(6) very well; d. Relation between the ratio $P_{+}/P_{-}$ to the ratio $\beta_{+}/\beta_{-}$. The numerical results are completely consistent with the theoretical prediction (red line) in Eq.(5).
Therefore, we can see from the above two examples of different potentials that our main predictions, i.e. Eq.(5) and Eq.(6), are very accurate and match the numerical results very well. We also need to stress that the relations in Eq.(5) and Eq.(6) are robust against by changing other parameters. In numerics, we already checked that our findings in Eq.(5) and Eq.(6) are robust against: (I) varying the temperatures $T$, (II) varying the coefficient $\Gamma$, (III) adding $\phi^{3}$ terms in the potential (7) and (IV) changing the coefficients in front of $\phi^{4}$. Therefore, it reflects that our predictions Eq.(5) and Eq.(6) depending only on $\sqrt{V^{\prime\prime}_{+}/V^{\prime\prime}_{-}}$ are very strong and universal predictions.
IV Conclusions and Discussions
@@
We studied the unequal probabilities of a field to roll down an asymmetric potential, in which the top is only $C^{1}$ continuous. By using the Fokker-Planck equation in stochastic equation, we theoretically found that the probability to roll down to the left (right) depends on the square root of the second derivative of the potential at that top, i.e. our main conclusions are the Eq.(5) and Eq.(6). Then we used two kinds of potentials in the time-dependent Ginzburg-Landau equation to numerically verify our conclusions. We found that our theoretical predictions are robust against varying other parameters. Therefore, our findings are universal.
There are many possible future uses for this new mechanism of asymmetric symmetry breaking. For instance, in cosmology one usually use the Higgs-type potential to study the cosmological phase transitions. If we change the potential to be like in our work, the probabilities to get the vacuum states will be different. This would change the cosmological phase transitions accordingly. Potentially, this new mechanism may play a role in the origins of asymmetries in the early Universe.
@@ Acknowledgement
—Appendix—
Appendix A Counting statistics of the binomial distribution
A.1 With Higgs-type potential
aRefer to captionb
Refer to captioncRefer to captiond
Figure S1: Counting statistics with Higgs-type potential in Eq.(7). a. The histogram of $P(n,k)$ for three different trial numbers $n$, with the probability $P_{+}=2/3$; b. The histogram of $P(n,k)$ for three different trial numbers $n$, with the probability $P_{+}=2-\sqrt{2}$; c. The vanishing probability of $P(n,k=0)$ in $n$ trials with $P_{+}=2/3$. The inset is the logarithmic plot, in which the slope of the straight line is $\log(1/3)$, consistent with the prediction $P_{+}=(1-P_{+})^{n}$; d. The vanishing probability of $P(n,k=0)$ in $n$ trials with $P_{+}=2-\sqrt{2}$. The inset is the logarithmic plot, in which the slope of the straight line is $\log(\sqrt{2}-1)$, consistent with the prediction $P_{+}=(1-P_{+})^{n}$.
In this appendix, we show the histograms of the probability distribution of rolling down to right. If we regard the rolling down to the right side as a success with probability $P_{+}$, then rolling down to the left is a failure with success $P_{-}=1-P_{+}$. Therefore, each time of the rolling down is a Bernoulli trial. Then after $n$ times of trials, the probability of rolling down to the right as $k$ times will satisfy the binomial distributions
$B(n,k)=\binom{n}{k}P_{+}^{k}(1-P_{+})^{n-k}.$ (S1)
In Fig.S1 we show the counting statistics of the rolling down to the right side with the Higgs-type potential in Eq.(7). For the left column, i.e., Fig.S1 a and c, we set $\alpha_{-}=1$ and $\alpha_{+}=4$. Therefore, the probability to roll down to the right side $P_{+}=\frac{\sqrt{4}}{1+\sqrt{4}}=2/3$ according to Eq.(6). While for the right column, i.e., Fig.S1 b and d, we set $\alpha_{-}=1$ and $\alpha_{+}=2$, thus $P_{+}=\frac{\sqrt{2}}{1+\sqrt{2}}=2-\sqrt{2}$. Panel a and b of Fig.S1 show the histogram of the probability $P(n,k)$, which means $k$ times of rolling down to the right side in $n$ trials. We can see that for the different probabilities $P_{+}$ and different trials numbers $n$, the histogram of $P(n,k)$ are consistent with the theoretical binomial distributions $B(n,k)$ in Eq.(S1) very well.
Fig.S1 c and d shows the vanishing probability to roll down to the right in $n$ trials, i.e., $P(n,k=0)$. Theoretically, from the binomial distribution Eq.(S1), this vanishing probability should be $B(n,k=0)=(1-P_{+})^{n}$. From the two plots we see that the numerical statistics match the theoretical predictions very well.
@@ A.1 With Higgs-type potential
A.2 With cosine-type potential
aRefer to captionb
Refer to captioncRefer to captiond
@@ A.2 With cosine-type potential
Figure S2: Counting statistics with cosine-type potential in Eq.(8). a. The histogram of $P(n,k)$ for three different trial numbers $n$, with the probability $P_{+}=3/5$; b. The histogram of $P(n,k)$ for three different trial numbers $n$, with the probability $P_{+}=3/4$; c. The vanishing probability of $P(n,k=0)$ in $n$ trials with $P_{+}=3/5$. The inset is the logarithmic plot, in which the slope of the straight line is $\log(2/5)$, consistent with the prediction $P_{+}=(1-P_{+})^{n}$; d. The vanishing probability of $P(n,k=0)$ in $n$ trials with $P_{+}=3/4$. The inset is the logarithmic plot, in which the slope of the straight line is $\log(1/4)$, consistent with the prediction $P_{+}=(1-P_{+})^{n}$.
@@
In Fig.S2 we show the counting statistics of the rolling down to the right side with the cosine-type potential in Eq.(8). For the left column, i.e., Fig.S2 a and c, we set $\beta_{-}=2$ and $\beta_{+}=3$. Therefore, the probability to roll down to the right side $P_{+}=\frac{3/2}{1+3/2}=3/5$ according to Eq.(6). While for the right column, i.e., Fig.S2 b and d, we set $\beta_{-}=2$ and $\beta_{+}=6$, thus $P_{+}=\frac{6/2}{1+6/2}=3/4$.
@@
From panel a and b of Fig.S2, we can see that for the different probabilities $P_{+}$ and different trials numbers $n$, the histogram of $P(n,k)$ are consistent with the theoretical binomial distributions $B(n,k)$ in Eq.(S1) very well.
@@
Fig.S2 c and d shows the vanishing probability to roll down to the right in $n$ trials, i.e., $P(n,k=0)$. Theoretically, from the binomial distribution Eq.(S1), this vanishing probability should be $B(n,k=0)=(1-P_{+})^{n}$. From the two plots we see that the numerical statistics match the theoretical predictions very well.
@@
Therefore, the consistency between the numerics and the binomial distributions in turn implies that our predictions of the $P_{+}$ in Eq.(6) are correct.

### v2

@@ II.1 Counting statistics of the binomial distribution
Figure 3: Counting statistics with Higgs-type potential in Eq.(7). a. Histogram of $P(n,k)$ for three different trial numbers $n$ with the probability $P_{+}=2/3$; b. Histogram of $P(n,k)$ for three different trial numbers $n$ with the probability $P_{+}=2-\sqrt{2}$; c. The probability of $P(n,k=0)$ in $n$ trials with $P_{+}=2/3$. The inset is the logarithmic plot, in which the slope of the straight line is $\log(1/3)$, consistent with the prediction $B(n,k=0)=(1-P_{+})^{n}$; d. The probability of $P(n,k=0)$ in $n$ trials with $P_{+}=2-\sqrt{2}$. The inset is the logarithmic plot, in which the slope of the straight line is $\log(\sqrt{2}-1)$, also matches the prediction $B(n,k=0)=(1-P_{+})^{n}$.
In Fig.3 we show the counting statistics of the rolling down to the right side with the Higgs-type potential. In the left column, i.e., Fig.3 a and c, we set $\alpha_{-}=1$ and $\alpha_{+}=4$. Therefore, according to our theoretical prediction in Eq.(6), the probability of rolling down to the right side is $P_{+}=\frac{\sqrt{4}}{1+\sqrt{4}}=2/3$. While in the right column, i.e., Fig.3 b and d, we set $\alpha_{-}=1$ and $\alpha_{+}=2$, thus $P_{+}=\frac{\sqrt{2}}{1+\sqrt{2}}=2-\sqrt{2}$. Panel a and b of Fig.3 show the histogram of the probability $P(n,k)$, which means the probability for $k$ times of rolling down to the right side in $n$ trials. We can see that for different probabilities $P_{+}$ and different trial numbers $n$, the histogram of $P(n,k)$ are consistent with the theoretical binomial distributions $B(n,k)$ in Eq.(12) very well. We did not show the error bars of the histogram since the errors are very tiny.
Fig.3 c and d exhibit the extremal cases that there are no events for rolling down to the right side in $n$ trials, i.e., $P(n,k=0)$, for different $P_{+}$. Theoretically, from the binomial distribution Eq.(12), this probability for vanishing $k$ should be $B(n,k=0)=(1-P_{+})^{n}$. From Fig.3 c and d, we see that the numerical statistics (circles) match the theoretical predictions (red lines) very well. The inset plots show the relations between $\log P(n,k=0)$ and $n$. From the binomial distribution, the theoretical relation should be $\log B(n,k=0)=n\log(1-P_{+})$, which are reflected by the linear relations in the inset plots. Moreover, the slopes are consistent with the theoretical predictions. Specifically, the slopes are $\log(1-P_{+})=\log(1-2/3)=\log(1/3)$ in Fig.3 c and $\log(1-P_{+})=\log(1-(2-\sqrt{2}))=\log(\sqrt{2}-1)$ in Fig.3 d, respectively.
Therefore, we see that the binomial distributions with the given probability $P_{+}$ are consistent with the numerical statistics. This in turn implies that our theoretical prediction in Eq.(5) and Eq.(6) are correct.
We must stress that our main findings in Eq.(5) and Eq.(6) are robust against by varying other parameters. In numerics, we already checked that they are robust against: (I) varying the reservoir temperature $T$; (II) varying the coefficient $\Gamma$; (III) adding $\phi^{3}$ terms in the potential Eq.(7); (IV) changing the coefficients in front of $\phi^{4}$; (V) adding higher order terms in the potential, such as $\phi^{5}$ and etc. In the Supplemental Materials we adopt another cosine-type potential to numerically verifies our theoretical predictions. Therefore, all of these reflect that our predictions in Eq.(5) and Eq.(6) are robust and universal.
III Conclusions and Discussions
@@
We studied the unequal probabilities of a field to roll down an asymmetric potential, in which the top is only $C^{1}$ continuous. By using the Fokker-Planck equation in stochastic process, we theoretically found that the probability to roll down to the left (right) depends on the square root of the second derivative of the potential at that top, i.e. our main conclusions are the Eq.(5) and Eq.(6). Then we used the Higgs-type potentials in the time-dependent Ginzburg-Landau equation to numerically verify our conclusions. We also found that our theoretical predictions are robust against varying other parameters. Therefore, our findings are universal results.
Our findings are important to the asymmetries of the vacuum selection. Potentially, this new mechanism may play an important role in explaining the origins of asymmetries in the early Universe given that one can build such an asymmetric effective potential first.
@@ Acknowledgement
Appendix A Unequal probabilities of vacuum selection with cosine-type potential
We will use another potential to verify the correctness of our main findings in Eq.(5) and Eq.(6) in the main text. It is the cosine-type potential,
$V(\phi)=\theta(-\phi)(\cos(\beta_{-}\phi)+\phi^{4})+\theta(\phi)(\cos(\beta_{+}\phi)+\phi^{4}).$ (S1)
In numerics, we fix $\beta_{-}\equiv 2$ and vary $\beta_{+}$ in the range of $\beta_{+}\in[0.2,8]$. Now we change the time step to be $\Delta t=0.02$ and still set $\Gamma=5$ and temperature $T=0.01$.
Panel a of Fig.S1 shows the profiles and the derivative of the potential (S1) with $\beta_{+}=6$. Obviously, the potential satisfies our requirement of the $\mathring{C}^{1}$ continuous. The minimum values of potential sits at $\phi_{-}=-0.7937$ and $\phi_{+}=0.5089$.
Panel b of Fig.S1 exhibits the time evolutions of the scalar fields. There are respectively three cases for each of them to approach $\phi_{+}$ (the upper lines) and $\phi_{-}$ (the lower lines) at equilibrium time. The inset plot shows the initial fluctuations of the scalar fields around the top, and then continuously roll down the potential as the arrows indicate.
Panel c of Fig.S1 is the relation between $P_{+}$ and the ratio $\beta_{+}/\beta_{-}$. From the potential (S1) we see that $\beta_{+}/\beta_{-}=\sqrt{V^{\prime\prime}_{+}/V^{\prime\prime}_{-}}$, therefore, the relations shown in panel c, i.e., $P_{+}=(\beta_{+}/\beta_{-})/(1+\beta_{+}/\beta_{-})$, exactly satisfies our prediction in Eq.(6) in the main text.
Panel d of Fig.S1 shows the relations between $P_{+}/P_{-}$ and $\beta_{+}/\beta_{-}$. The linear relation $P_{+}/P_{-}=\beta_{+}/\beta_{-}$ in numerics and theoretical predictions Eq.(5) in the main text match each other very well.
aRefer to captionbRefer to captioncRefer to captiond
Figure S1: a. The cosine-type potential $V(\phi)$ (blue) in Eq.(S1) with $\beta_{-}=2$ and $\beta_{+}=6$, and its derivative $V^{\prime}(\phi)$ (red); b. Time time evolution of the scalar fields. Each has three cases to approach positive and negative minimums. The inset plot shows the very early fluctuations around the top due to the temperature fluctuation. After some time, they will persistently roll down the potential as the arrows indicate; c. Relation between the probability $P_{+}$ to the ratio $\beta_{+}/\beta_{-}$. Numerical results match the theoretical prediction (red line) Eq.(6) in the main text very well; d. Relation between the ratio $P_{+}/P_{-}$ to the ratio $\beta_{+}/\beta_{-}$. The numerical results are completely consistent with the theoretical prediction (red line) in Eq.(5) in the main text.
Appendix B Binomial distribution with cosine-type potential
aRefer to captionbRefer to captioncRefer to captiond
@@ Appendix B Binomial distribution with cosine-type potential
Figure S2: Counting statistics with cosine-type potential in Eq.(S1). a. The histogram of $P(n,k)$ for three different trial numbers $n$, with the probability $P_{+}=3/5$; b. The histogram of $P(n,k)$ for three different trial numbers $n$, with the probability $P_{+}=3/4$; c. The vanishing probability of $P(n,k=0)$ in $n$ trials with $P_{+}=3/5$. The inset is the logarithmic plot, in which the slope of the straight line is $\log(2/5)$, consistent with the prediction $B(n,k=0)=(1-P_{+})^{n}$; d. The vanishing probability of $P(n,k=0)$ in $n$ trials with $P_{+}=3/4$. The inset is the logarithmic plot, in which the slope of the straight line is $\log(1/4)$, consistent with the prediction $B(n,k=0)=(1-P_{+})^{n}$.
@@
In Fig.S2 we show the counting statistics of the rolling down to the right side with the cosine-type potential in Eq.(S1). For the left column, i.e., Fig.S2 a and c, we set $\beta_{-}=2$ and $\beta_{+}=3$. Therefore, the probability to roll down to the right side is $P_{+}=\frac{3/2}{1+3/2}=3/5$ according to Eq.(6) in the main text. While for the right column, i.e., Fig.S2 b and d, we set $\beta_{-}=2$ and $\beta_{+}=6$, thus $P_{+}=\frac{6/2}{1+6/2}=3/4$.
@@
From panel a and b of Fig.S2, we can see that for the different probabilities $P_{+}$ and different trials numbers $n$, the histogram of $P(n,k)$ are consistent with the theoretical binomial distributions $B(n,k)$ in Eq.(12) in the main text very well.
@@
Fig.S2 c and d shows the probability to roll down to the right with zero times in $n$ trials, i.e., $P(n,k=0)$. Theoretically, from the binomial distribution Eq.(12) in the main text, this probability should be $B(n,k=0)=(1-P_{+})^{n}$. The two inset plots show the corresponding logarithmic figures, i.e., $\log P(n,k=0)$ vs. $n$. Specifically, from the inset plots we can see that the slopes are exactly the value of $\log(1-P_{+})$ since $\log B(n,k=0)=n\log(1-P_{+})$. Therefore, from the two plots we see that the numerical statistics match the theoretical predictions very well.
@@
Consequently, we find that by using the cosine-type potential (S1), the consistency between the numerics and the binomial distributions in turn implies that our main predictions in Eq.(5) and Eq.(6) in the main text are correct.